# Compositional Generative Inverse Design

### 24 Jan 2024 v1

@@ Abstract
Inverse design, where we seek to design input variables in order to optimize an underlying objective function, is an important problem that arises across fields such as mechanical engineering to aerospace engineering. Inverse design is typically formulated as an optimization problem, with recent works leveraging optimization across learned dynamics models. However, as models are optimized they tend to fall into adversarial modes, preventing effective sampling. We illustrate that by instead optimizing over the learned energy function captured by the diffusion model, we can avoid such adversarial examples and significantly improve design performance. We further illustrate how such a design system is compositional, enabling us to combine multiple different diffusion models representing subcomponents of our desired system to design systems with every specified component. In an N-body interaction task and a challenging 2D multi-airfoil design task, we demonstrate that by composing the learned diffusion model at test time, our method allows us to design initial states and boundary shapes that are more complex than those in the training data. Our method outperforms state-of-the-art neural inverse design method by an average of 41.5% in prediction MAE and 14.3% in design objective for the N-body dataset and discovers formation flying to minimize drag in the multi-airfoil design task. Project website and code can be found at https://github.com/AI4Science-WestlakeU/cindm.
@@ 1 Introduction
Concretely, we contribute the following: (1) We propose a novel formulation for inverse design as an energy optimization problem. (2) We introduce Compositional Inverse Design with Diffusion Models (CinDM) method, which enables us to generalize to out-of-distribution and more complex design inputs than seen in training. (3) We present a set of benchmarks for inverse design in 1D and 2D. Our method outperforms prior state-of-the-art models by an average of 41.5% in prediction MAE and 14.3% in design objective for the N-body dataset and discovers formation flying to minimize drag in the multi-airfoil design task.
@@ 3.2 Generative Inverse Design
1:  Require Compositional set of diffusion models $\epsilon^{i}_{\theta}(z_{s},s),i=1,2,...N$, design objective $\mathcal{J}(\cdot)$, covariance matrix $\sigma_{t}^{2}I$, hyperparameters $\lambda,S,K$
@@ 3.3 Compositional Generative Inverse Design
II. Generalization to more interacting bodies. Many inverse design applications require generalizing the trained model to more interacting bodies for a dynamical system. Our method allows such generalization by composing the energy function of few-body interactions to more interacting bodies. Now we illustrate it with a 2-body to N-body generalization. Suppose that only the trajectory of a 2-body interaction is given, where we have the trajectory of $U^{(i)}_{[0,T]}=(U^{(i)}_{0},U^{(i)}_{1},...,U^{(i)}_{T})$ for body $i\in\{1,2\}$. We can learn an energy function $E_{\theta}((U^{(1)}_{[0,T]},U^{(2)}_{[0,T]}),\gamma)$ from this trajectory. In the test time, given $N>2$ interacting bodies subjecting to the same pairwise interactions, the energy function for the combined trajectory $U_{[0,T]}=(U^{(1)}_{[0,T]},...,U^{(N)}_{[0,T]})$ for the $N$ bodies is then given by:
@@ 4 Experiments
In the experiments, we aim to answer the following questions: (1) Can CinDM generalize to more complex designs in the test time using its composition capability? (2) Comparing backpropagation with surrogate models and other strong baselines, can CinDM improve on the design objective? (3) Can CinDM address high-dimensional design space? To answer these questions, we perform our experiments in three different scenarios: compositional inverse design in time dimension (Sec. 4.1), compositional inverse design generalizing to more objects (Sec. 4.2), and 2D compositional design for multiple airfoils with Navier-Stokes flow (Sec. 4.3). Each of the above experiments represents an important scenario in inverse design and has important implications in science and engineering. In each experiment, we compare CinDM with the state-of-the-art deep learning-based inverse design method proposed by Allen et al. (2022), which we term Backprop, and cross-entropy method (CEM) (Rubinstein & Kroese, 2004) which is a standard sampling-based optimization method typically used in classical inverse design. Additionally, we compare CinDM with two inverse-design methods: neural adjoint method with the boundary loss function (NABL) and conditional invertible neural network (cINN) method (Ren et al., 2020; Ansari et al., 2022). The details and results are provided in Appendix H, in which we adopt a more reasonable experimental setting to those new baselines. All the baselines and our model contain similar numbers of parameters in each comparison for fair evaluation. To evaluate the performance of each inverse design method, we feed the output of the inverse design method (i.e., the optimized initial or boundary states) to the ground-truth solver, perform rollout by the solver and feed the rollout trajectory to the design objective. We do not use the surrogate model to perform rollout since the trained surrogate models may not be faithful to the ground-truth dynamics and can overestimate the design objective. By evaluating using a ground-truth solver, all inverse design methods can be evaluated fairly.
@@ 4.1 Compositional inverse design in time
In this experiment, we aim to test each method’s ability to generalize to more forward time steps than during training. This is important since in test time, the inverse design methods are typically used over longer predictions horizons than in training. We use an N-body interaction environment where each ball with a radius of 0.1 is bouncing in a $1\times 1$ box. The balls will exchange momentum when elastically colliding with each other or with the wall,. The design task is to identify the initial state (position and velocity of the balls) of the system such that the end state optimizes a certain objective (e.g., as close to a certain target as possible). This setting represents a simplified version of many real-life scenarios such as billiard, bowling, and ice hockey. Since the collisions preserve kinetic energy but modify speed and direction of each ball and multiple collisions can happen over a long time, this represents a non-trivial inverse design problem with abrupt changes in the design space. During training time, we provide each method with training trajectory consisting of 24 steps, and in test time, let it roll out for a total of 24, 34, 44, and 54 steps. The design objective is to minimize the last step’s Euclidean distance to the center $(x,y)=(0.5,0.5)$. For baselines, we compare with CEM (Rubinstein & Kroese, 2004) and Backprop (Allen et al., 2022). Each method uses either Graph Network Simulator (GNS, Sanchez-Gonzalez et al. (2020), a state-of-the-art method for modeling N-body interactions) or U-Net (Ronneberger et al., 2015) as backbone architecture that either predicts 1 step or 23 steps in a single forward pass. For our method, we use the same U-Net backbone architecture for diffusion. To perform time composition, we superimpose $N$ EBMs $E_{\theta}(U_{[0,T]},\gamma)$ on states with overlapping time ranges: $U_{[0,23]}$, $U_{[10,33]}$, $U_{[20,43]}$,… $U_{[10(N-1),10(N-1)+23]}$ as in Eq. 7, and use Eq. 10 to perform denoising diffusion. Besides evaluating with the design objective ($\mathcal{J}$), we also use the metric of mean absolute error (MAE) between the predicted trajectory and the trajectory generated by the ground-truth solver to evaluate how faithful each method’s prediction is. Each design scenario is run 50 times and the average performance is reported in Table 1. We show example trajectories of our method in Fig. 2. Details for the architecture and training are provided in Appendix A. We also make comparison with a simple baseline that performs diffusion over 44 steps directly without time composition. Details and results are presented in Appendix C.
@@
CEM, GNS (1-step) 0.3021 0.14941 0.2531 0.13296 0.2781 0.20109 0.2845 0.19811
CEM, GNS 0.3144 0.12741 0.3178 0.16538 0.3102 0.24884 0.3059 0.24863
CEM, U-Net (1-step) 0.2733 0.08013 0.2680 0.13183 0.2910 0.14783 0.2919 0.13348
CEM, U-Net 0.2731 0.02995 0.2424 0.02937 0.2616 0.04460 0.2804 0.06520
Backprop, GNS (1-step) 0.1216 0.03678 0.1643 0.02976 0.1966 0.03645 0.2657 0.10331
Backprop, GNS 0.2453 0.13024 0.2822 0.11200 0.2959 0.12867 0.2877 0.14241
Backprop, U-Net (1-step) 0.2020 0.06338 0.2193 0.07705 0.2187 0.05668 0.2851 0.07716
Backprop, U-Net 0.1168 0.01137 0.1294 0.01303 0.1481 0.00804 0.3140 0.01675
From Table 1, we see that our method generally outperforms the baselines by a wide margin. In the “2-body 24 steps” scenario which is the same setting as in training and without composition, our model performs similarly to “Backprop with U-Net” that shares the same backbone architecture. However, with more prediction steps, the design objective and MAE for all the baseline methods worsen quickly. In contrast, our method’s metrics remain much lower than the baselines, with wider gaps for longer prediction steps. For example, our method’s design objective outperforms the best baseline by 2.1%, 3.3%, 10.5%, and 42.3% for 24, 34, 44, and 54-step predictions, respectively, with an average of 14.6% improvement. Similarly, our method’s MAE outperforms the best baseline by -5.7%, 41.4%, 13.6%, and 48.1% for 24, 34, 44, and 54-step predictions, respectively, with an average of 24.3% improvement. This shows the two-fold advantage of our method. Firstly, even with the same backbone architecture, our diffusion method can roll out stably and accurately for much longer than the baseline, since the forward surrogate models in the baselines during design may encounter out-of-distribution and adversarial inputs which it does not know how to evolve properly. On the other hand, our diffusion-based method is trained to denoise and favor inputs consistent with the underlying physics. Secondly, our compositional method allows our model to generalize to longer time steps and allows for stable rollout. An example trajectory designed by our CinDM is shown in Fig. 2 (a). We see that it matches with the ground-truth simulation nicely, captures the bouncing with walls and with other balls, and the end position of the bodies tends towards the center, showing the effectiveness of our method. We also see that Backprop’s performance are superior to the sampling-based CEM, consistent with Allen et al. (2022).
@@ 4.2 Compositional inverse design generalizing to more objects
In this experiment, we test each method’s performance in inverse design on larger state dimensions than in training. We utilize the N-body simulation environment as in Sec. 4.1, but instead of considering longer trajectories, we test on more bodies than in training. This setting is also inspired by real-life scenarios where the dynamics in test time have more interacting objects than in training (e.g., in astronomical simulation and biophysics). Specifically, all methods are trained with only 2-body interactions with 24 time steps, and tested with 4-body and 8-body interactions for 24 and 44 time steps using Eq. 8. This is a challenging task since the methods need to generalize to a much larger state space than in training. For N-body interaction, there are $N(N-1)/{2}$ pairs of 2-body interactions. The case with 44 time steps adds difficulty by testing generalization in both state size and time (composing $28\times 3=84$ diffusion models for CinDM).
@@
CEM, GNS (1-step) 0.3029 0.20027 0.3215 0.26518 0.3312 0.36865 0.3292 0.37430
CEM, GNS 0.3139 0.21253 0.3110 0.26924 0.3221 0.26708 0.3319 0.32678
Backprop, GNS (1-step) 0.2872 0.08023 0.2900 0.11331 0.3312 0.27988 0.3227 0.74314
Backprop, GNS 0.3118 0.10249 0.3423 0.15277 0.3302 0.19039 0.3233 0.24718
CinDM (ours) 0.2066 0.04152 0.2281 0.03195 0.3056 0.08821 0.3169 0.09566
From Table 2, we see that our CinDM method outperforms all baselines by a wide margin in both the design objective and MAE. On average, our method achieves an improvement of 14.1% in design objective, and an improvement of 58.8% in MAE than the best baseline. In Fig. 2 (b), we see that our method captures the interaction of the 4 bodies with the wall and each other nicely and all bodies tend towards center at the end. The above results again demonstrate the strong compositional capability of our method: it can generalize to much larger state space than seen in training.
@@ Appendix B Full results for compositional inverse design of the N-body task
Here we provide the full statistical results including the 95% confidence interval (for 50 instances) for N-body experiments, including compostional inverse design in time and more objects. Specifically, Table 6 shows detailed results for Table 1 in Section 4.1; and Table 7 extends Table 2 in Section 4.2.
@@
CEM, GNS (1-step) 0.3021 ± 0.0266 0.14941 ± 0.02772 0.2531 ± 0.0212 0.13296 ± 0.02692 0.2781 ± 0.0227 0.20109 ± 0.03369 0.2845 ± 0.0211 0.19811 ± 0.03721
CEM, GNS 0.3144 ± 0.0302 0.12741 ± 0.02432 0.3178 ± 0.0215 0.16538 ± 0.02322 0.3102 ± 0.0198 0.24884 ± 0.02011 0.3059 ± 0.0190 0.24863 ± 0.01410
CEM, U-Net (1-step) 0.2733 ± 0.0268 0.08013 ± 0.02803 0.2680 ± 0.0207 0.13183 ± 0.03146 0.2910 ± 0.0207 0.14783 ± 0.03025 0.2919 ± 0.0242 0.13348 ± 0.03039
CEM, U-Net 0.2731 ± 0.0260 0.02995 ± 0.00921 0.2424 ± 0.0176 0.02937 ± 0.00889 0.2616 ± 0.0199 0.04460 ± 0.01406 0.2804 ± 0.0247 0.06520 ± 0.01932
Backprop, GNS (1-step) 0.1216 ± 0.0053 0.03678 ± 0.00666 0.1643 ± 0.0186 0.02976 ± 0.00449 0.1966 ± 0.0252 0.03645 ± 0.00404 0.2657 ± 0.0222 0.10331 ± 0.04124
Backprop, GNS 0.2453 ± 0.0242 0.13024 ± 0.02470 0.2822 ± 0.0191 0.11200 ± 0.01560 0.2959 ± 0.0196 0.12867 ± 0.01608 0.2877 ± 0.0186 0.14241 ± 0.01569
Backprop, U-Net (1-step) 0.2020 ± 0.0158 0.06338 ± 0.00905 0.2193 ± 0.0226 0.07705 ± 0.02047 0.2187 ± 0.0196 0.05668 ± 0.01464 0.2851 ± 0.0219 0.07716 ± 0.02250
Backprop, U-Net 0.1168 ± 0.0042 0.01137 ± 0.00202 0.1294 ± 0.0092 0.01303 ± 0.00207 0.1481 ± 0.0110 0.00804 ± 0.00164 0.3140 ± 0.0137 0.01675 ± 0.00188
CinDM (ours) 0.1143 ± 0.0047 0.01202 ± 0.00114 0.1251 ± 0.0071 0.00763 ± 0.00069 0.1326 ± 0.0087 0.00695 ± 0.00067 0.1533 ± 0.0140 0.00870 ± 0.00150
@@
CEM, GNS (1-step) 0.3029 ± 0.0129 0.20027 ± 0.02875 0.3215 ± 0.0143 0.26518 ± 0.02648 0.3312 ± 0.0084 0.36865 ± 0.02384 0.3292 ± 0.0062 0.37430 ± 0.02231
CEM, GNS 0.3139 ± 0.0127 0.21253 ± 0.02030 0.3110 ± 0.0147 0.26924 ± 0.01362 0.3221 ± 0.0075 0.26708 ± 0.01043 0.3319 ± 0.0078 0.32678 ± 0.00991
Backprop, GNS (1-step) 0.2872 ± 0.0114 0.08023 ± 0.01550 0.2900 ± 0.0170 0.11331 ± 0.02922 0.3312 ± 0.0068 0.27988 ± 0.02224 0.3227 ± 0.0063 0.74314 ± 0.02940
Backprop, GNS 0.3118 ± 0.0136 0.10249 ± 0.01311 0.3423 ± 0.0141 0.15277 ± 0.01279 0.3302 ± 0.0063 0.19039 ± 0.00797 0.3233 ± 0.0061 0.24718 ± 0.00705
CinDM (ours) 0.2066 ± 0.0118 0.04152 ± 0.00644 0.2281 ± 0.0145 0.03195 ± 0.00705 0.3056 ± 0.0062 0.08821 ± 0.00593 0.3169 ± 0.0068 0.09566 ± 0.00924

### 11 Mar 2024 v2

@@ Abstract
Inverse design, where we seek to design input variables in order to optimize an underlying objective function, is an important problem that arises across fields such as mechanical engineering to aerospace engineering. Inverse design is typically formulated as an optimization problem, with recent works leveraging optimization across learned dynamics models. However, as models are optimized they tend to fall into adversarial modes, preventing effective sampling. We illustrate that by instead optimizing over the learned energy function captured by the diffusion model, we can avoid such adversarial examples and significantly improve design performance. We further illustrate how such a design system is compositional, enabling us to combine multiple different diffusion models representing subcomponents of our desired system to design systems with every specified component. In an N-body interaction task and a challenging 2D multi-airfoil design task, we demonstrate that by composing the learned diffusion model at test time, our method allows us to design initial states and boundary shapes that are more complex than those in the training data. Our method generalizes to more objects for N-body dataset and discovers formation flying to minimize drag in the multi-airfoil design task. Project website and code can be found at https://github.com/AI4Science-WestlakeU/cindm.
@@ 1 Introduction
Concretely, we contribute the following: (1) We propose a novel formulation for inverse design as an energy optimization problem. (2) We introduce Compositional Inverse Design with Diffusion Models (CinDM) method, which enables us to generalize to out-of-distribution and more complex design inputs than seen in training. (3) We present a set of benchmarks for inverse design in 1D and 2D. Our method generalizes to more objects for N-body dataset and discovers formation flying to minimize drag in the multi-airfoil design task.
@@ 3.2 Generative Inverse Design
1:  Require Compositional set of diffusion models $\epsilon^{i}_{\theta}(z_{s},s),i=1,2,...N$, design objective $\mathcal{J}(\cdot)$, covariance matrix $\sigma_{s}^{2}I$, hyperparameters $\lambda,S,K$
@@ 3.3 Compositional Generative Inverse Design
II. Generalization to more interacting bodies. Many inverse design applications require generalizing the trained model to more interacting bodies for a dynamical system, which is far more difficult than generalizing to more time steps. Our method allows such generalization by composing the energy function of few-body interactions to more interacting bodies. Now we illustrate it with a 2-body to N-body generalization. Suppose that only the trajectory of a 2-body interaction is given, where we have the trajectory of $U^{(i)}_{[0,T]}=(U^{(i)}_{0},U^{(i)}_{1},...,U^{(i)}_{T})$ for body $i\in\{1,2\}$. We can learn an energy function $E_{\theta}((U^{(1)}_{[0,T]},U^{(2)}_{[0,T]}),\gamma)$ from this trajectory. In the test time, given $N>2$ interacting bodies subjecting to the same pairwise interactions, the energy function for the combined trajectory $U_{[0,T]}=(U^{(1)}_{[0,T]},...,U^{(N)}_{[0,T]})$ for the $N$ bodies is then given by:
@@ 4 Experiments
In the experiments, we aim to answer the following questions: (1) Can CinDM generalize to more complex designs in the test time using its composition capability? (2) Comparing backpropagation with surrogate models and other strong baselines, can CinDM improve on the design objective or prediction accuracy? (3) Can CinDM address high-dimensional design space? To answer these questions, we perform our experiments in three different scenarios: compositional inverse design in time dimension (Sec. 4.1), compositional inverse design generalizing to more objects (Sec. 4.2), and 2D compositional design for multiple airfoils with Navier-Stokes flow (Sec. 4.3). Each of the above experiments represents an important scenario in inverse design and has important implications in science and engineering. In each experiment, we compare CinDM with the state-of-the-art deep learning-based inverse design method proposed by Allen et al. (2022), which we term Backprop, and cross-entropy method (CEM) (Rubinstein & Kroese, 2004) which is a standard sampling-based optimization method typically used in classical inverse design. Additionally, we compare CinDM with two inverse-design methods: neural adjoint method with the boundary loss function (NABL) and conditional invertible neural network (cINN) method (Ren et al., 2020; Ansari et al., 2022). The details and results are provided in Appendix H, in which we adopt a more reasonable experimental setting to those new baselines. All the baselines and our model contain similar numbers of parameters in each comparison for fair evaluation. To evaluate the performance of each inverse design method, we feed the output of the inverse design method (i.e., the optimized initial or boundary states) to the ground-truth solver, perform rollout by the solver and feed the rollout trajectory to the design objective. We do not use the surrogate model to perform rollout since the trained surrogate models may not be faithful to the ground-truth dynamics and can overestimate the design objective. By evaluating using a ground-truth solver, all inverse design methods can be evaluated fairly.
@@ 4.1 Compositional inverse design in time
In this experiment, we aim to test each method’s ability to generalize to more forward time steps than during training. This is important since in test time, the inverse design methods are typically used over longer predictions horizons than in training. We use an N-body interaction environment where each ball with a radius of 0.1 is bouncing in a $1\times 1$ box. The balls will exchange momentum when elastically colliding with each other or with the wall. The design task is to identify the initial state (position and velocity of the balls) of the system such that the end state optimizes a certain objective (e.g., as close to a certain target as possible). This setting represents a simplified version of many real-life scenarios such as billiard, bowling, and ice hockey. Since the collisions preserve kinetic energy but modify speed and direction of each ball and multiple collisions can happen over a long time, this represents a non-trivial inverse design problem with abrupt changes in the design space. During training time, we provide each method with training trajectory consisting of 24 steps, and in test time, let it roll out for a total of 24, 34, 44, and 54 steps. The design objective is to minimize the last step’s Euclidean distance to the center $(x,y)=(0.5,0.5)$. For baselines, we compare with CEM (Rubinstein & Kroese, 2004) and Backprop (Allen et al., 2022). Each method uses either Graph Network Simulator (GNS, Sanchez-Gonzalez et al. (2020), a state-of-the-art method for modeling N-body interactions) or U-Net (Ronneberger et al., 2015) as backbone architecture that either predicts 1 step or 23 steps in a single forward pass. For our method, we use the same U-Net backbone architecture for diffusion. To perform time composition, we superimpose $N$ EBMs $E_{\theta}(U_{[0,T]},\gamma)$ on states with overlapping time ranges: $U_{[0,23]}$, $U_{[10,33]}$, $U_{[20,43]}$,… $U_{[10(N-1),10(N-1)+23]}$ as in Eq. 7, and use Eq. 10 to perform denoising diffusion. Besides evaluating with the design objective ($\mathcal{J}$), we also use the metric of mean absolute error (MAE) between the predicted trajectory and the trajectory generated by the ground-truth solver to evaluate how faithful each method’s prediction is. Each design scenario is run 500 times and the average performance is reported in Table 1. We show example trajectories of our method in Fig. 2. Details for the architecture and training are provided in Appendix A. We also make comparison with a simple baseline that performs diffusion over 44 steps directly without time composition. Details and results are presented in Appendix C.
@@
CEM, GNS (1-step) 0.2622 0.13963 0.2204 0.15378 0.2701 0.21277 0.2773 0.21706
CEM, GNS 0.2699 0.12746 0.3142 0.14637 0.3056 0.18155 0.3124 0.20266
CEM, U-Net (1-step) 0.2364 0.07720 0.2391 0.09701 0.2744 0.11885 0.2729 0.12992
CEM, U-Net 0.1762 0.03597 0.1639 0.03094 0.1816 0.03900 0.1887 0.04350
Backprop, GNS (1-step) 0.1452 0.04339 0.1497 0.03806 0.1511 0.03621 0.1851 0.04104
Backprop, GNS 0.2407 0.09788 0.2678 0.11017 0.2762 0.12395 0.2952 0.13963
Backprop, U-Net (1-step) 0.2182 0.07554 0.2445 0.08278 0.2536 0.08487 0.2751 0.10599
Backprop, U-Net 0.1228 0.01974 0.1171 0.01236 0.1143 0.00970 0.1289 0.01067
CinDM (ours) 0.1160 0.01264 0.1288 0.00917 0.1447 0.00959 0.1650 0.01064
From Table 1, we see that our method is competitive in design objectives and outperforms every baseline in MAE. In the “2-body 24 steps” scenario which is the same setting as in training and without composition, our method outperforms the strongest baselines by a wide margin both on design objective and MAE. With more prediction steps, our method not only performs better than any baselines in MAE but also merely is weaker than the strongest baseline in design objective. For example, our method’s MAE outperforms the best baseline by 36.0%, 25.8%, 1.1%, and 0.3% for 24, 34, 44, and 54-step predictions, respectively, with an average of 15.8% improvement. Similarly, our method’s design objective outperforms the best baseline by 5.5% for 24-step. This shows the two-fold advantage of our method. Firstly, even with the same backbone architecture, our diffusion method can roll out stably and accurately for much longer than the baseline, since the forward surrogate models in the baselines during design may encounter out-of-distribution and adversarial inputs which it does not know how to evolve properly. On the other hand, our diffusion-based method is trained to denoise and favor inputs consistent with the underlying physics. Secondly, our compositional method allows our model to generalize to longer time steps and allows for stable rollout. An example trajectory designed by our CinDM is shown in Fig. 2 (a). We see that it matches with the ground-truth simulation nicely, captures the bouncing with walls and with other balls, and the end position of the bodies tends towards the center, showing the effectiveness of our method. We also see that Backprop’s performance are superior to the sampling-based CEM, consistent with Allen et al. (2022).
@@ 4.2 Compositional inverse design generalizing to more objects
In this experiment, we test each method’s performance in inverse design on larger state dimensions than in training. We utilize the N-body simulation environment as in Sec. 4.1, but instead of considering longer trajectories, we test on more bodies than in training. This setting is also inspired by real-life scenarios where the dynamics in test time have more interacting objects than in training (e.g., in astronomical simulation and biophysics). Specifically, all methods are trained with only 2-body interactions with 24 time steps, and tested with 4-body and 8-body interactions for 24 and 44 time steps using Eq. 8. This is a markedly more challenging task than generalizing to more time steps since the methods need to generalize to a much larger state space than in training. For N-body interaction, there are $N(N-1)/{2}$ pairs of 2-body interactions. The case with 44 time steps adds difficulty by testing generalization in both state size and time (composing $28\times 3=84$ diffusion models for CinDM).
@@
CEM, GNS (1-step) 0.3173 0.23293 0.3307 0.53521 0.3323 0.38632 0.3306 0.53839
CEM, GNS 0.3314 0.25325 0.3313 0.28375 0.3314 0.25325 0.3313 0.28375
Backprop, GNS (1-step) 0.2947 0.06008 0.2933 0.30416 0.3280 0.46541 0.3317 0.72814
Backprop, GNS 0.3221 0.09871 0.3195 0.15745 0.3251 0.15917 0.3299 0.21489
CinDM (ours) 0.2034 0.03928 0.2254 0.03163 0.3062 0.09241 0.3212 0.09249
From Table 2, we see that our CinDM method outperforms all baselines by a wide margin in both the design objective and MAE. On average, our method achieves an improvement of 15.6% in design objective, and an improvement of 53.4% in MAE than the best baseline. In Fig. 2 (b), we see that our method captures the interaction of the 4 bodies with the wall and each other nicely and all bodies tend towards center at the end. The above results again demonstrate the strong compositional capability of our method: it can generalize to much larger state space than seen in training.
@@ Appendix B Full results for compositional inverse design of the N-body task
Here we provide the full statistical results including the 95% confidence interval (for 500 instances) for N-body experiments, including compostional inverse design in time and more objects. Specifically, Table 6 shows detailed results for Table 1 in Section 4.1; and Table 7 extends Table 2 in Section 4.2.
@@
CEM, GNS (1-step) 0.2622 ± 0.0090 0.13963 ± 0.00999 0.2204 ± 0.0072 0.15378 ± 0.01123 0.2701 ± 0.0079 0.21277 ± 0.01264 0.2773 ± 0.0070 0.21706 ± 0.01256
CEM, GNS 0.2699 ± 0.0081 0.12746 ± 0.00662 0.3142 ± 0.0064 0.14637 ± 0.00657 0.3056 ± 0.0060 0.18155 ± 0.00689 0.3124 ± 0.0062 0.20266 ± 0.00679
CEM, U-Net (1-step) 0.2364 ± 0.0068 0.07720 ± 0.00623 0.2391 ± 0.0081 0.09701 ± 0.00796 0.2744 ± 0.0073 0.11885 ± 0.00854 0.2729 ± 0.0074 0.12992 ± 0.00897
CEM, U-Net 0.1762 ± 0.0071 0.03597 ± 0.00395 0.1639 ± 0.0062 0.03094 ± 0.00342 0.1816 ± 0.0072 $0$ 0.00451 0.1887 ± 0.0075 0.04350 ± 0.00487
Backprop, GNS (1-step) 0.1452 ± 0.0050 0.04339 ± 0.00285 0.1497 ± 0.0061 0.03806 ± 0.00304 0.1511 ± 0.0062 0.03621 ± 0.00322 0.1851 ± 0.0062 0.04104 ± 0.00285
Backprop, GNS 0.2407 ± 0.0067 0.09788 ± 0.00615 0.2678 ± 0.0072 0.11017 ± 0.00620 0.2762 ± 0.0071 0.12395 ± 0.00657 0.2952 ± 0.0073 0.13963 ± 0.00623
Backprop, U-Net (1-step) 0.2182 ± 0.0068 0.07554 ± 0.00466 0.2445 ± 0.0093 0.08278 ± 0.00613 0.2536 ± 0.0078 0.08487 ± 0.00611 0.2751 ± 0.0088 0.10599 ± 0.00709
Backprop, U-Net 0.1228 ± 0.0040 0.01974 ± 0.00223 0.1171 ± 0.0032 0.01236 ± 0.00104 0.1143 ± 0.0026 0.00970 ± 0.00076 0.1289 ± 0.0043 0.01067 ± 0.00090
CinDM (ours) 0.1160 ± 0.0019 0.01264 ± 0.00057 0.1288 ± 0.0030 0.00917 ± 0.00070 0.1447 ± 0.0040 0.00959 ± 0.00116 0.1650 ± 0.0045 0.01064 ± 0.00117
@@
CEM, GNS (1-step) 0.3173 ± 0.0040 0.23293 ± 0.01007 0.3307 ± 0.0022 0.53521 ± 0.00987 0.3323 ± 0.0023 0.38632 ± 0.00737 0.3306 ± 0.0023 0.53839 ± 0.01001
CEM, GNS 0.3314 ± 0.0023 0.25325 ± 0.00369 0.3313 ± 0.0023 0.28375 ± 0.00336 0.3314 ± 0.0023 0.25325 ± 0.00369 0.3313 ± 0.0023 0.28375 ± 0.00336
Backprop, GNS (1-step) 0.2947 ± 0.0044 0.06008 ± 0.00437 0.2933 ± 0.0041 0.30416 ± 0.03387 0.3280 ± 0.0026 0.46541 ± 0.02768 0.3317 ± 0.0023 0.72814 ± 0.01783
Backprop, GNS 0.3221 ± 0.0043 0.09871 ± 0.00499 0.3195 ± 0.0042 0.15745 ± 0.00561 0.3251 ± 0.0021 0.15917 ± 0.00261 0.3299 ± 0.0022 0.21489 ± 0.00238
CinDM (ours) 0.2034 ± 0.0032 0.03928 ± 0.00161 0.2254 ± 0.0044 0.03163 ± 0.00251 0.3062 ± 0.0021 0.09241 ± 0.00210 0.3212 ± 0.0023 0.09249 ± 0.00276